# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

### v3

@@ IV-D2 The Improvement on Personalization Performance
This section explores the impact of integrating the Talk2Drive framework into autonomous driving systems to enhance personalization based on the LLM model GPT-4 [openai_gpt-4_2023]. One of the central focuses of our system is its ability to offer drivers a personalized driving experience. Through Talk2Drive, individuals can express their preferences or feelings with varying degrees of directness, prompting the system to make corresponding adjustments. The commands collected are divided into three levels of directness, as explained in Tab. V.
@@
For each command directness within each expected driving behavior, the system’s responses to various responses given by human participants are collected and calculated into the defined metrics. Here, we utilize the takeover rate as our primary metric, and we assess the takeover rate both pre- and post-integration of Talk2Drive. We incorporate a rule-based system with a keyword-trigger logic as the baseline for comparison. For instance, in the highway scenario, commands like “accelerate,” “left,” and “right” prompt the system to execute relevant actions such as speeding up or changing lanes accordingly. Results are gathered in diverse driving scenarios on human drivers with diverse driving styles.
TABLE VII: Takeover Rate Improvement through Personalization.
@@
As shown in Tab. VII, the takeover rate in all driving scenarios decreases significantly, ranging from a 56.3% reduction to the complete elimination of takeover cases, demonstrating the Talk2Drive system’s ability to personalize the driving experience based on human preference. Additionally, one important finding from our results is that for all levels of command directness, our Talk2Drive system shows significant improvements compared to the baseline systems.
IV-D3 The Effectiveness of the Memory Module
To further investigate the performance of our memory module, we also conduct experiments comparing personalization across two settings: conditions without the memory module, and conditions with the memory module. These experiments take place in the parking lot, where the vehicle follows a pre-defined trajectory and adjusts speed based on human inputs. Humans take over in this context when they find the adjusted speed unsatisfactory. As demonstrated in Tab. VIII, there are significant reductions in takeover rate with Talk2Drive, regardless of the driver’s aggression or conservatism levels.
The results in Tab. VIII reveal that the inclusion of the memory module leads to a marked reduction in the takeover rate. The utilization of Talk2Drive framework without the memory module brings takeover rates between 0.14 and 0.29. When implementing Talk2Drive framework with the memory module, the takeover rate can be further decreased to only 0.07. Compared to their personalization performance, Talk2Drive with a memory module can reduce takeover rates by up to 65.2% compared to those without the memory module, which illustrates the benefits of a history-recording module in achieving a more personalized driving experience.
TABLE VIII: Effectiveness of Memory Modules in Influencing Takeover Rates
Driver Without Memory Module With Memory Module
A 0.14 0.07 (50.0$\downarrow$)
B 0.23 0.08 (65.2$\downarrow$)
C 0.29 0.18 (37.9$\downarrow$)
V Real-Vehicle Experiments: On-Board VLM for Personalized Motion Control
V-A Problem Statement
We adopt the traditional module-based autonomous driving framework that includes the full pipeline from perception to motion control, and our focus is specifically on enhancing the decision-making process at the motion control level, adapting the control performance to accommodate different human driving styles. The goal of the proposed system is to translate both verbal commands $\boldsymbol{I}$ and visual inputs $\boldsymbol{V}$ into executable control sequences for the motion control process. The onboard VLM acts as a translation mechanism $f:(\boldsymbol{I},\boldsymbol{V})\rightarrow\boldsymbol{P}$ that generates a policy $\boldsymbol{P}$, which is then fed into predefined maneuver programs.
Additionally, system messages $\boldsymbol{S}$ are sent to our VLM to specify both the tasks and adjustment strategies. In our setting, similar to the Sec. IV, $\boldsymbol{S}$ is generated through a predefined language generator.
Simultaneously, to further enhance personalization, we implement a RAG-based memory module to build a database storing historical human-vehicle interactions. Whenever a human activates this system, only relevant historical scenarios $\boldsymbol{H}$ are retrieved and provided to the VLM as reference. After each trip, users can provide feedback $\boldsymbol{F}$ on the generated control policy $\boldsymbol{P}$ for the current situations (including instructions $\boldsymbol{I}$ and visual input $\boldsymbol{V}$), which helps refine the VLM’s reasoning process. Subsequently, the instructions $\boldsymbol{I}$, scene description $\boldsymbol{D}$ (generated by on-board VLM), policy $\boldsymbol{P}$, and feedback $\boldsymbol{F}$ are packaged as a historical vector-based data entry and stored in the RAG database. Therefore, there are three procedures in our system:
@@ V-A Problem Statement
V-B Our Framework
V-B1 System Input
@@
As illustrated in Fig. 9, our fine-tuned on-board VLM processes four critical inputs for reasoning. The primary inputs consist of visual data $\boldsymbol{V}$ from the onboard camera and natural language commands $\boldsymbol{I}$ which are converted from verbal human instructions $\boldsymbol{I}$ using the open-source local speech recognition model “Whisper [radford2022robustspeechrecognitionlargescale].” Notably, our system leverages contextual and environmental information captured in the visual inputs $\boldsymbol{V}$, including weather conditions, traffic conditions, and road characteristics. For instance, the system automatically adopts a more conservative driving policy during nighttime operations or adverse weather conditions.
@@
Additionally, a pre-defined system message generator is employed to produce a customized system message $\boldsymbol{S}$, which is then simultaneously sent to the VLM.
@@
Furthermore, the VLM incorporates relevant interaction history $\boldsymbol{H}$ extracted from our RAG-based memory module as contextual input, which includes previous human instructions $\boldsymbol{I}$, scene descriptions $\boldsymbol{D}$, executed actions $\boldsymbol{A}$, and user feedback $\boldsymbol{F}$. Note that the scene descriptions $\boldsymbol{D}$ are generated by the on-board VLM to describe the image $\boldsymbol{V}$. This historical context enables the VLM to generate more appropriate responses by considering past interactions and human feedback. A detailed discussion of how our memory module works will be presented in subsection V-B3.
V-B2 Reasoning and Action Generation
@@
In our approach, reasoning within the VLM framework enables the interpretation of diverse driving scenarios and user instructions to generate actionable outputs. Traditional controllers in motion control typically rely on a default set of parameters; however, following the approach in [sha2023languagempclargelanguagemodels], our VLM will generate two distinct action matrices to separately manage the PID controller for longitudinal movement and the MPC for lateral movement. These matrices translate the model’s understanding of the environment and user preferences into precise control actions, guiding the autonomous vehicle’s behavior. Specifically, they are used by the controllers to generate acceleration $a$ and steering angle $\delta_{f}$, which are executed by the vehicle’s ECU. The ECU then sends low-level control signals to the drive-by-wire system developed by AutonomousStuff [autostuff], enabling smooth and responsive vehicle operation. The general process of this subsection can be shown below:
Output Action $P:=\begin{bmatrix}K_{p}&K_{i}&K_{d}\\ W_{l}&W_{h}&W_{s}\end{bmatrix}$ (62)
@@
V-B3 RAG-Enhanced Memory Module
@@
$\{I-(I,D,P,F)\}$ (63)
When processing a new driving scenario, the instruction $\boldsymbol{I}$ is used for similarity matching to retrieve the top-k similar prior situations. The associated data values are then sent to the VLM, enhancing decision-making with relevant context and supporting personalization. This RAG-enhanced memory enables the VLM to handle similar situations with greater accuracy and consistency, improving the vehicle’s responsiveness to unique user preferences and enhancing the overall driving experience.
V-B4 Multi-Controller Joint Motion Control
@@
$J=E^{T}QE+\Delta_{f}^{T}R\Delta_{f}$ (65)
where $E$ is the predicted state calculated by the prediction model, and $\Delta_{f}$ is the future control input. The $Q$ and $R$ are weighting matrices that penalize tracking state deviations and control effort. Our VLM generates the action matrix that primarily considers three key components: the weight for lateral error ($W_{l}\in Q$), the weight for heading error ($W_{h}\in Q$), and the weight for the squared terms of speed and steering inputs ($W_{s}\in R$). These weighting factors are selected as they demonstrate the most significant impact on lateral control performance.
V-B5 Efficient On-Board VLM Module
@@
The human instructions are also very diverse, ranging from explicit commands like “speed up” to more implicit cues such as “I am in an urgent situation.” This diversity allows the VLM to interpret both clear and vague inputs, improving its ability to understand complex human intentions. To enhance the model’s responsiveness to different driving styles, we annotate each image with three different instruction types—aggressive, moderate, and conservative—each paired with a corresponding action. This multi-faceted approach ensures that the VLM can adapt its behavior to match various driving styles, enabling it to respond flexibly and contextually across diverse traffic conditions.
@@
We apply the LoRA method to fine-tune our Qwen-VL model. LoRA works by freezing the pre-trained model weights and introducing trainable, low-rank decomposition matrices into each layer of the Transformer architecture. This approach significantly reduces the number of trainable parameters required, making fine-tuning more efficient.
The fine-tuning process for our VLM is conducted on a cluster of four NVIDIA A100 GPUs, each equipped with 40GB of memory. The model is trained over five epochs with a per-device batch size of two for training and one for validation, using a learning rate of 1e-5. Additionally, we implemented gradient accumulation with eight steps, allowing for effective larger batch processing. This setup enables the entire training process to be completed in approximately five hours, ensuring both accuracy and efficiency in model performance.
@@
V-C Real-World Experiment
To comprehensively evaluate our system’s performance, we conduct a series of experiments assessing its ability to provide safe, comfortable, reliable, and personalized driving experiences. We employ multiple evaluation metrics: a driving score to measure driving performance, including safety, comfort, and alignment with environmental conditions and human instructions; takeover frequency to assess personalization capabilities. Additionally, we perform an ablation study to examine the effectiveness of the memory module. The demo for our real-vehicle experiment is publicly available [^3] Real-vehicle experiment demo video with on-board VLM: https://youtu.be/lxzXhTasSdMover.
TABLE IX: Driving Performance Validation. $\downarrow$: Lower Values are Better. $\uparrow$: Higher Values are Better.
@@ V-C Real-World Experiment
Collision($s$$\uparrow$
$SV_{x}$
($m^{2}/s^{2}$$\downarrow$
$SV_{y}$
($10^{-2}m^{2}/s^{2}$$\downarrow$
@@
Lane Change Baseline 2.44 3.91 1.65 0.37 3.14 0.88 1.01 - 88.5 60.0 74.5
@@
Left Turn Baseline - 1.12 7.52 0.22 1.81 1.29 1.36 - 88.0 60.0 70.4
@@
V-C1 Experiment Setup
Test Track and Participants
The field experiments [^4] The experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. aim at validating the real-world performance of our personalized motion control system. We include three types of driving behaviors—accelerating, lane changing, and turning—to comprehensively validate control over steering, throttle, and braking. An overview of the test track and driving behaviors is shown in Fig. 10. For both acceleration and lane change scenarios, a lead vehicle is positioned 30 $m$ ahead of the ego vehicle, accelerating from static to 45 $km/h$ with an acceleration of 1.26 $m/s^{2}$. In the acceleration scenario, the ego vehicle accelerates from a complete stop to reach 50 $km/h$. In the lane change scenario, the ego vehicle maintains 50 $km/h$ while overtaking the lead vehicle. For the intersection turning scenario, the ego vehicle navigates a curve with a radius of 23.89 $m$ at a constant speed of 30 $km/h$.
V-C2 System Driving Performance
@@
The models are then assessed based on four key aspects—safety, comfort, time efficiency, and alignment, as shown in Tab. IX. An overall driving score $S$ is then calculated as a weighted sum of the individual metric scores, denoted as:
$S=\sum{w_{k}\cdot S_{k}}$ (66)
where $k$ includes all ten metrics: Time to Collision ($\tau$), longitudinal and lateral speed variance ($SV_{x}$ and $SV_{y}$), lateral and longitudinal mean absolute acceleration ($|\bar{\alpha_{x}}|$ and $|\bar{\alpha_{y}}|$), lateral and longitudinal mean absolute jerk ($|\bar{J_{x}}|$ and $|\bar{J_{y}}|$), VLM Latency, Command Alignment and Scenario Alignment. All the scores $S_{k}$ range from 0 to 100, while the weights of scores $w_{k}$ are empirically tuned for each driving scenario. For instance, longitudinal parameters have higher weights in acceleration scenarios, while lateral parameters are weighted more heavily in turning scenarios. The safety and comfort metrics are followed as defined in Sec. IV-C3.
@@
$S_{K_{p}}^{*}=\begin{cases}\frac{100(K_{p}-K_{p,\min})}{K_{p,\text{lower}}-K_{p,\min}},&K_{p}\in[K_{p,\min},K_{p,\text{lower}}),\\[10.0pt] 100,&K_{p}\in[K_{p,\text{lower}},K_{p,\text{upper}}),\\[10.0pt] \frac{100(K_{p,\max}-K_{p})}{K_{p,\max}-K_{p,\text{upper}}},&K_{p}\in[K_{p,\text{upper}},K_{p,\max}],\\[10.0pt] 0,&\text{else}.\end{cases}$
@@
As shown in Tab. IX, the command alignment score of our VLM model is similar to or greater than GPT-4o, showing our model has high reasoning capabilities regarding the personalization of control parameters. As for the scenario alignment, our model shows significantly better performance than baseline conditions in lane changing and left turn scenarios but a score very similar to the baseline scenario. We think this is mostly due to the model considering acceleration on a clear straight road in less dangerous situations and thus does not act too conservatively. In terms of the overall Driving Score, our model also surpasses the performance of baseline models and even GPT-4o in some scenarios, indicating our model provides the most well-rounded driving experience.
V-C3 Human-in-the-Loop Validation

### v4

@@ V-D2 The Improvement on Personalization Performance
This section explores the impact of integrating the Talk2Drive framework into autonomous driving systems to enhance personalization based on the LLM model GPT-4 [openai_gpt-4_2023]. One of the central focuses of our system is its ability to offer drivers a personalized driving experience. Through Talk2Drive, individuals can express their preferences or feelings with varying degrees of directness, prompting the system to make corresponding adjustments. The commands collected are divided into three levels of directness, as explained in Tab. VII.
@@
For each command directness within each expected driving behavior, the system’s responses to various responses given by human participants are collected and calculated into the defined metrics. Here, we utilize the takeover rate as our primary metric and evaluate it both pre- and post-integration of Talk2Drive. We incorporate a rule-based system with a keyword-trigger logic as the baseline for comparison. For instance, in the highway scenario, commands like “accelerate,” “left,” and “right” prompt the system to execute relevant actions such as speeding up or changing lanes accordingly. The results are gathered in diverse driving scenarios on human drivers with diverse driving styles.
TABLE IX: Takeover Rate Improvement through Personalization.
@@
As shown in Tab. IX, the takeover rate in all driving scenarios decreases significantly, ranging from a 56.3% reduction to the complete elimination of takeover cases, demonstrating the Talk2Drive system’s ability to personalize the driving experience based on human preference. Additionally, one important finding from our results is that for all levels of command directness, our Talk2Drive system shows significant improvements compared to the baseline systems.
V-D3 The Effectiveness of the Memory Module
To further investigate the performance of our memory module, we also conduct experiments comparing personalization across two settings: conditions without the memory module, and conditions with the memory module. These experiments take place in the parking lot, where the vehicle follows a pre-defined trajectory and adjusts speed based on human inputs. Humans take over in this context when they find the adjusted speed unsatisfactory. As demonstrated in Tab. X, there are significant reductions in takeover rate with Talk2Drive, regardless of the driver’s aggression or conservatism levels.
The results in Tab. X reveal that the inclusion of the memory module leads to a marked reduction in the takeover rate. The utilization of Talk2Drive framework without the memory module brings takeover rates between 0.14 and 0.29. When implementing Talk2Drive framework with the memory module, the takeover rate can be further decreased to only 0.07. Compared to their personalization performance, Talk2Drive with a memory module can reduce takeover rates by up to 65.2% compared to those without the memory module, which illustrates the benefits of a history-recording module in achieving a more personalized driving experience.
TABLE X: Effectiveness of Memory Modules in Influencing Takeover Rates
Driver
Without Memory Module
With Memory Module
@@ V-D3 The Effectiveness of the Memory Module
VI Real-Vehicle Experiments: On-Board VLM for Personalized Motion Control
VI-A Problem Statement
We adopt the traditional module-based autonomous driving framework that includes the full pipeline from perception to motion control, and our focus is specifically on enhancing the decision-making process at the motion control level, adapting the control performance to accommodate different human driving styles. The goal of the proposed system is to translate both verbal commands $I$ and visual inputs $V$ into executable control sequences for the motion control process. The on-board VLM acts as a translation mechanism $f:(I,V)\rightarrow P$ that generates a policy $P$, which is then fed into predefined maneuver programs.
Additionally, system messages $S$ are sent to our VLM to specify both the tasks and adjustment strategies. In our setting, we follow the Sec. V, $S$ is generated through a predefined language generator.
Simultaneously, to further enhance personalization, we implement a RAG-based memory module to build a database storing historical human-vehicle interactions. Whenever a human activates this system, only relevant historical scenarios $H$ are retrieved and provided to the VLM as reference. After each trip, users can provide feedback $F$ on the generated control policy $P$ for the current situations (including instructions $I$ and visual input $V$), which helps refine the VLM’s reasoning process. Subsequently, the instructions $I$, scene description $C$ (generated by on-board VLM), policy $P$, and feedback $F$ are packaged as a historical vector-based data entry and stored in the RAG database. Therefore, there are three procedures in our system:
$\text{VLM Execution}:$ $P\xleftarrow{\text{VLM}}f(I,S,V,H);$ (63)
$\text{Human Feedback}:$ $F\xleftarrow{\text{Human}}[I,V,P];$
$\text{Memory Update}:$ $H\leftarrow\left[I,C,P,F\right]$
VI-B Our Framework
VI-B1 System Input
@@
As illustrated in Fig. 11, our fine-tuned on-board VLM processes four critical inputs for reasoning. The primary inputs consist of visual data $V$ from the on-board camera and natural language commands $I$ which are converted from verbal human instructions $I$ using the open-source local speech recognition model “Whisper [radford2022robustspeechrecognitionlargescale].” Notably, our system leverages contextual and environmental information captured in the visual inputs $V$, including weather conditions, traffic conditions, and road characteristics. For instance, the system automatically adopts a more conservative driving policy during nighttime operations or adverse weather conditions.
@@
Additionally, a pre-defined system message generator is employed to produce a customized system message $S$, which is then simultaneously sent to the VLM.
@@
Furthermore, the VLM incorporates relevant interaction history $H$ extracted from our RAG-based memory module as contextual input, which includes previous human instructions $I$, scene descriptions $C$, executed actions $A$, and user feedback $F$. Note that the scene descriptions $C$ are generated by the on-board VLM to describe the image $V$. This historical context enables the VLM to generate more appropriate responses by considering past interactions and human feedback. A detailed discussion of how our memory module works will be presented in subsection VI-B3.
VI-B2 Reasoning and Action Generation
@@
In our approach, reasoning within the VLM framework enables the interpretation of diverse driving scenarios and user instructions to generate actionable outputs. Traditional controllers in motion control typically rely on a default set of parameters; however, following the approach in [sha2023languagempclargelanguagemodels], our VLM will generate two distinct action matrices to separately manage the PID controller for longitudinal movement and the model predictive control (MPC) for lateral movement. These matrices translate the model’s understanding of the environment and user preferences into precise control actions, guiding the autonomous vehicle’s behavior. Specifically, they are used by the controllers to generate acceleration $a$ and steering angle $\delta_{f}$, which are executed by the vehicle’s ECU. The ECU then sends low-level control signals to the drive-by-wire system developed by AutonomousStuff [autostuff], enabling smooth and responsive vehicle operation. The general process of this subsection can be shown below:
Output Action $P=\begin{bmatrix}K_{p}&K_{i}&K_{d}\\ W_{l}&W_{h}&W_{s}\end{bmatrix}$ (64)
@@
VI-B3 RAG-Enhanced Memory Module
@@
$\{I-(I,C,P,F)\}$ (65)
When processing a new driving scenario, the instruction $I$ is used for similarity matching to retrieve the top-k similar prior situations. The associated data values are then sent to the VLM, enhancing decision-making with relevant context and supporting personalization. This RAG-enhanced memory enables the VLM to handle similar situations with greater accuracy and consistency, improving the vehicle’s responsiveness to unique user preferences and enhancing the overall driving experience.
VI-B4 Multi-Controller Joint Motion Control
@@
$J=E^{T}W_{Q}E+\Delta_{f}^{T}W_{R}\Delta_{f}$ (67)
where $E$ is the predicted state calculated by the prediction model, and $\Delta_{f}$ is the future control input. The $W_{Q}$ and $W_{R}$ are weighting matrices that penalize tracking state deviations and control effort. Our VLM generates the action matrix that primarily considers three key components: the weight for lateral error ($W_{l}\in W_{Q}$), the weight for heading error ($W_{h}\in W_{Q}$), and the weight for the squared terms of speed and steering inputs ($W_{s}\in W_{R}$). These weighting factors are selected as they demonstrate the most significant impact on lateral control performance.
VI-B5 Efficient On-Board VLM Module
@@
The human instructions are also very diverse, ranging from explicit commands like “speed up” to more implicit cues such as “I am in an urgent situation.” This diversity allows the VLM to interpret both clear and vague inputs, improving its ability to understand complex human intentions. To enhance the model’s responsiveness to different driving styles, we annotate each image with three different instruction types (aggressive, moderate, and conservative), each paired with a corresponding action. This approach ensures that the VLM can adapt its behavior to match various driving styles, enabling it to respond flexibly and contextually across diverse traffic conditions.
@@
We apply the LoRA method to fine-tune our Qwen-VL model. LoRA operates by freezing the pre-trained model weights and introducing trainable, low-rank decomposition matrices into each layer of the Transformer architecture. This approach significantly reduces the number of trainable parameters required, making fine-tuning more efficient.
The fine-tuning process of our VLM is conducted on a cluster of four NVIDIA A100 GPUs, each equipped with 40GB of memory. The model is trained over five epochs with a per-device batch size of two for training and one for validation, using a learning rate of 1e-5. Additionally, we implemented gradient accumulation with eight steps, allowing for effective larger batch processing. This setup enables the entire training process to be completed in approximately five hours, ensuring both accuracy and efficiency in model performance.
@@
VI-C Real-World Experiment
To comprehensively evaluate our system’s performance, we conduct a series of experiments assessing its ability to provide safe, comfortable, reliable, and personalized driving experiences. We employ multiple evaluation metrics: a driving score to measure driving performance, including safety, comfort, and alignment with environmental conditions and human instructions; takeover frequency to assess personalization capabilities. Additionally, we perform an ablation study to examine the effectiveness of the memory module. The demo for our real-vehicle experiment is publicly available [^3] Real-vehicle experiment demo video with on-board VLM: https://www.youtube.com/watch?v=IPZo2USvkAw.
TABLE XI: Driving Performance Validation. $\downarrow$: Lower Values are Better. $\uparrow$: Higher Values are Better.
Driving
Scenario
@@ VI-C Real-World Experiment
[…]
$\sigma_{x}$
[…]
$\sigma_{y}$
[…]
@@
Lane
Change
Baseline 2.44 3.91 1.65 0.37 3.14 0.88 1.01 - 88.5 60.0 74.5
@@
Left
Turn
Baseline - 1.12 7.52 0.22 1.81 1.29 1.36 - 88.0 60.0 70.4
@@
VI-C1 Experiment Setup
The field experiments [^4] All experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. aim at validating the real-world performance of our personalized motion control system. We include three types of driving behaviors (accelerating, lane changing, and turning) to comprehensively validate control over steering, throttle, and braking. An overview of the test track and driving behaviors is shown in Fig. 12. For both acceleration and lane change scenarios, a lead vehicle is positioned 30 $m$ ahead of the ego vehicle, accelerating from static to 45 $km/h$ with an acceleration of 1.26 $m/s^{2}$. In the acceleration scenario, the ego vehicle accelerates from a complete stop to reach 50 $km/h$. In the lane change scenario, the ego vehicle maintains 50 $km/h$ while overtaking the lead vehicle. For the intersection turning scenario, the ego vehicle navigates a curve with a radius of 23.89 $m$ at a constant speed of 30 $km/h$. This study includes seven participants with diverse demographic characteristics. The participants consisted of 61.4% male and 28.6% female drivers, with ages ranging from 23 to 30 years (mean = 26.42, std = 3.24). Their driving experience varies considerably (mean = 6.42 years, std = 4.27 years). All participants hold valid U.S. driving licenses.
VI-C2 System Driving Performance
@@
The models are then assessed based on four key aspects: safety, comfort, time efficiency, and alignment, as shown in Tab. XI. An overall driving score $Score$ is then calculated as a weighted sum of the individual metric scores, denoted as:
$Score=\sum{w_{k}\cdot Score_{k}}$ (68)
where $k$ includes all ten metrics: Time to Collision ($\tau$), longitudinal and lateral speed variance ($\sigma_{x}$ and $\sigma_{y}$), lateral and longitudinal mean absolute acceleration ($|\bar{\alpha_{x}}|$ and $|\bar{\alpha_{y}}|$), lateral and longitudinal mean absolute jerk ($|\bar{J_{x}}|$ and $|\bar{J_{y}}|$), VLM Latency, Command Alignment and Scenario Alignment. All the scores $Score_{k}$ range from 0 to 100, while the weights of scores $w_{k}$ are empirically tuned for each driving scenario. For instance, longitudinal parameters have higher weights in acceleration scenarios, while lateral parameters are weighted more heavily in turning scenarios. The safety and comfort metrics are followed as defined in Sec. V-C3.
@@
$Score_{K_{p}}^{*}=\begin{cases}\frac{100(K_{p}-K_{p,\min})}{K_{p,\text{lower}}-K_{p,\min}},&K_{p}\in[K_{p,\min},K_{p,\text{lower}}),\\[10.0pt] 100,&K_{p}\in[K_{p,\text{lower}},K_{p,\text{upper}}),\\[10.0pt] \frac{100(K_{p,\max}-K_{p})}{K_{p,\max}-K_{p,\text{upper}}},&K_{p}\in[K_{p,\text{upper}},K_{p,\max}],\\[10.0pt] 0,&\text{else}.\end{cases}$
@@
As shown in Tab. XI, the command alignment score of our VLM model is similar to or greater than GPT-4o, showing our model has high reasoning capabilities regarding the personalization of control parameters. As for the scenario alignment, our model shows significantly better performance than baseline conditions in lane changing and left turn scenarios but a score very similar to the baseline scenario. We think this is mostly due to the model considering acceleration on a clear straight road in less dangerous situations and thus does not act too conservatively. In terms of the overall Driving Score, our model also surpasses the performance of baseline models and even GPT-4o in some scenarios, indicating our model provides the most well-rounded driving experience.
VI-C3 Human-in-the-Loop Validation